The following shows a list of weak acids and bases with their Ka and Kb values. Substances Ka Substances Kb Cyanic acid (HCN) 4.9 x 10-10 Ammonia (NH3) 1.8 x 10-5 Nitrous acid (HNO2) 4.5 x 10-4 Aniline (C6H5NH2) 4.3 x 10-10 Hydrofluoric acid (HF) 6.8 x 10-4 Pyridine (C5H5N) 1.7 x 10-9 Acetic acid (CH3COOH) 1.8 x 10-5 Methylamine (CH3NH2) 4.4 x 10-4 If you want to make a pH 8.0 buffer solution, which of the following acids or bases will give you the best buffer solution (having greatest buffer capacity in both direction)? Suggest your method to prepare pH 8.0 buffer using your choice.
Answer #1


acids: greatest acid buffer capacity range pH = pKa+1 to pH = pka-1

Cyanic acid (HCN)   pKa = -log(4.9 x 10^(-10)) = 9.3

Nitrous acid (HNO2) pKa = -log(4.5 x 10^(-4)) = 3.347

Hydrofluoric acid (HF) pKa = -log (6.8 *10^(-4)) = 3.16

Acetic acid (CH3COOH) pKa = -log (1.8 x 10-5) = 4.74

bases : Basic buffer capacity range is maxium pH = 14-( pKb+1) to pH = 14-(pkb-1).

Ammonia (NH3) pKb= -log(1.8*10^(-5)) = 4.74

Aniline (C6H5NH2) pKb = -log(4.3* 10^(-10)) = 9.36

Pyridine (C5H5N) pKb = -log(1.7* 10^(-9)) = 8.77

Methylamine (CH3NH2) pKb = -log(4.4* 10^(-4)) = 3.356

from the above data most suitable one to prepare buffer solution with pH= 8 is

Ammonia (NH3) ,Because for basic buffer buffer capacity is maximum at pOH = pKb ,so that pH = 14-pOH

therefore NH3 pKb = 4.74 ,pH = 14-4.74 = 9.26

so take NH3 and add strong acid in 10:1 mole ratio you will get desire buffer.